# Mohamed Bouadi, Pratinav Seth Aditya Tanna, Vinay Kumar Sankarapu Lexsi Labs, India & France

### v1

@@ A.1  Pretraining Data Generation
Following the pretraining paradigm of TabICL [32], we train our model on synthetically generated datasets to learn generalizable representations for in-context learning on tabular data. Unlike natural language or vision domains where large-scale real data is available, tabular datasets exhibit extreme heterogeneity in schemas, distributions, and task objectives. Synthetic data generation via structural causal models (SCMs) enables us to control dataset diversity while ensuring coverage of diverse statistical patterns.
@@ A.1.1  Structural Causal Model (SCM) Prior
We generate synthetic datasets using SCM-based priors [3, 31], where features are related through nonlinear causal relationships. For a dataset with $m$ features, we define a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each node $v\in\mathcal{V}$ represents a feature, and edges $\mathcal{E}$ encode causal dependencies.
@@
  • Standard nonlinearities: ReLU, ReLU6 [23], SELU [22], SiLU (Swish) [15], Softplus
@@
    with $N=256$, $b_{i}\sim\mathcal{U}[0,2\pi]$, $a_{i}\sim\mathcal{U}[0,N]$, and $w_{i}:=a_{i}^{-\exp(u)}$ where $u\sim\mathcal{U}[0.7,3.0]$. This random function approximates complex, non-parametric relationships [33].
@@ A.1.2  Tree-Based SCM Prior
To complement the continuous SCM prior, we also generate datasets using a tree-based SCM prior [13], where the causal mechanism $f_{j}$ is a decision tree or random forest. This prior is particularly important for modeling categorical interactions and hierarchical decision boundaries commonly observed in real-world tabular data (e.g., credit scoring, medical diagnosis).
@@ A.2  Pretraining Details
Across all stages, we use the Adam optimizer [21] with gradient norm clipping at 1.0 and a learning rate schedule shown in Figure 3. This curriculum, progressing from small, uniform datasets to large, variable datasets with selective fine-tuning, enables the model to generalize effectively across diverse dataset scales while preventing overfitting to specific sample counts.
@@ A.3.1  Column-wise Embedding ($\text{TF}_{\text{col}}$)
A.3.2  Multi-Scale Sparse Row Interaction ($\text{TF}_{\text{row}}^{\text{MS}}$)
@@ A.3.3  Cross-Component Perceiver Memory
A.3.4  Dataset-wise In-Context Learning ($\text{TF}_{\text{icl}}$)
@@
A.4  Datasets

### v3

@@ A.1  Pretraining Data Generation
Following the pretraining paradigm of TabICL [tabicl], we train our model on synthetically generated datasets to learn generalizable representations for in-context learning on tabular data. Unlike natural language or vision domains where large-scale real data is available, tabular datasets exhibit extreme heterogeneity in schemas, distributions, and task objectives. Synthetic data generation via structural causal models (SCMs) enables us to control dataset diversity while ensuring coverage of diverse statistical patterns.
@@ A.1.1  Structural Causal Model (SCM) Prior
We generate synthetic datasets using SCM-based priors [scm, muller2022pfn], where features are related through nonlinear causal relationships. For a dataset with $m$ features, we define a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{V},\mathcal{E})$ where each node $v\in\mathcal{V}$ represents a feature, and edges $\mathcal{E}$ encode causal dependencies.
@@
  • Standard nonlinearities: ReLU, ReLU6 [relu6], SELU [selu], SiLU (Swish) [silu], Softplus
@@
    with $N=256$, $b_{i}\sim\mathcal{U}[0,2\pi]$, $a_{i}\sim\mathcal{U}[0,N]$, and $w_{i}:=a_{i}^{-\exp(u)}$ where $u\sim\mathcal{U}[0.7,3.0]$. This random function approximates complex, non-parametric relationships [rahimi2007random].
@@ A.1.2  Tree-Based SCM Prior
To complement the continuous SCM prior, we also generate datasets using a tree-based SCM prior [grinsztajn2022tree], where the causal mechanism $f_{j}$ is a decision tree or random forest. This prior is particularly important for modeling categorical interactions and hierarchical decision boundaries commonly observed in real-world tabular data (e.g., credit scoring, medical diagnosis).
@@ A.2  Pretraining Details
Across all stages, we use the Adam optimizer [kingma2014adam] with gradient norm clipping at 1.0 and a learning rate schedule shown in Figure 3. This curriculum, progressing from small, uniform datasets to large, variable datasets with selective fine-tuning, enables the model to generalize effectively across diverse dataset scales while preventing overfitting to specific sample counts.
@@ A.3.1  Column-wise Embedding ($\text{TF}_{\text{col}}$)
A.3.2  Multi-Scale Sparse Row Interaction
@@ A.3.3  Cross-Component Perceiver Memory
A.3.4  Dataset-wise In-Context Learning
@@
Appendix B Further Experiments
Figure 4 presents the relative accuracy improvement (%) of each method over XGBoost, evaluated per dataset across the three benchmarks: TALENT (Figure 4(a)), TabZilla (Figure 4(b)), and OpenML-cc18 (Figure 4(c)). Each point corresponds to a per-dataset delta, while boxplots summarize the distribution (median, interquartile range, and whiskers). The dashed vertical line denotes parity with XGBoost (0%).
Across all three benchmarks, Orion-MSP consistently improves upon the strong XGBoost baseline. On TabZilla and OpenML-cc18, Orion-MSP achieves positive median relative accuracy with a compact interquartile range and few negative outliers, indicating both higher accuracy and greater reliability. On TALENT, Orion-MSP reaches parity in median performance but exhibits lower variance than most neural baselines.
Among classical boosted trees, LightGBM, CatBoost, and Random Forest cluster tightly around parity, showing comparable behavior. In contrast, tabular foundation models (TFMs), notably Mitra and ContextTab, exhibit pronounced negative shifts and high variance. TabPFN and TabICL perform competitively and occasionally outperform Orion-MSP on specific datasets, yet their broader variance and heavier left tails reveal less consistent behavior. Overall, Orion-MSP matches or surpasses their central performance and achieves the best mean–variance trade-off, confirming the benefits of our model design for tabular generalization.
(a) Relative accuracy improvement over XGBoost on TALENT Benchmark
(b) Relative accuracy improvement over XGBoost on TabZilla Benchmark
(c) Relative accuracy improvement over XGBoost on OPENML-CC18 Benchmark
Figure 4: Relative accuracy improvement over XGBoost on three benchmarks.
To further quantify cross-dataset performance, we computed per-dataset ranks across all 11 methods (lower is better) for each benchmark, averaged them over datasets, and conducted a one-way Friedman test to assess overall differences. When significant, Nemenyi post-hoc tests were applied, and the resulting critical difference (CD) diagrams were plotted at $\alpha=0.05$; methods connected by the CD bar are not significantly different.
(a) Relative accuracy improvement over XGBoost on TALENT Benchmark
(b) Relative accuracy improvement over XGBoost on TabZilla Benchmark
(c) Relative accuracy improvement over XGBoost on OPENML-CC18 Benchmark
Figure 5:
As shown in Figure 5, Orion-MSP attains the best average rank on all three benchmarks—3.09 on TALENT, 3.32 on TabZilla, and 3.35 on OpenML-cc18—appearing as the leftmost method in each CD diagram. TabICL, TabPFN, and TabDPT follow closely, typically lagging by $\approx$ 0.5–1.5 rank points. Tree ensembles (XGBoost, LightGBM, CatBoost, Random Forest) occupy a middle tier with average ranks of 4.6–6.2, showing parity among themselves but a consistent gap to Orion-MSP and the stronger TFMs. Finally, ContextTab and Mitra form the rightmost group (ranks $\approx$ 9–10), confirming the underperformance seen in the improvement plots.
In summary, the CD diagrams corroborate our main finding: Orion-MSP is the top performer across diverse tabular benchmarks, outperforming tree ensembles on average and matching or exceeding pretrained tabular foundation models while maintaining a favorable significance profile.
Appendix C Datasets